Calculate ΔS when one mole of steam at 100 ℃ is converted into ice at 0 ℃. Average specific heat of water = 4.184 x 103 JK-1 kg-1; heat of vaporization at boiling point = 2490.6 x 103 J kg-1; heat of fusion at freezing point = 333.555 x 103 J kg-1.
1
Expert's answer
2020-12-06T13:51:50-0500

"\Delta"S = "\Delta"condensationS + "\Delta"coolingS + "\Delta"freezingS = - "\Delta"vaporizationH/Tvaporization + "\Delta"coolingS +

- "\Delta"meltingH/Tmelting = ((18g/mol)/(1000 g*kg-1))*(-(2490.6*103 J*kg-1)/(373.15 K) +

+ "\intop"373.15K273.15K[(4.184*103 J*K-1*kg-1)/T]dT - (333.555*103 J*kg-1)/(273.15 K)) = ((18g/mol)/(1000 g*kg-1))*(-7895.6699 + 4184*ln(273.15/373.15) J*K-1*kg-1) = -165.616 J*K-1*mol-1

We have 1 mole of water so "\Delta"S = (1 mol)*(-165.616 J*K-1*mol-1) = -165.616 J*K-1